Assume Aduhene economy does not trade with the outside world. In a given fiscal year, households spent 80% of their disposable income on consumption in addition to 600 consumption expenditure which is independent of income. Total government expenditure of 900 was supposed to be financed from a proportion tax levy of 40% on national income and a lump sum tax of 450. Total private investment spending is 700. Given that aggregate spending is equal to aggregate output.
1. Determine the value of the multiplier
2. Find the equilibrium level of output
1
Expert's answer
2020-04-27T07:43:53-0400

1."Multiplier = 1 / (\text{sum of the propensity to save} + tax + import)"

tax=0.4

"MPS=\frac{\text{change in savings}}{\text{change in income}}" =480/600=0.8

multiplier=1/(0.8+0.4)=0.83

2.At equilibrium "AD=AS"

"output=Consumption + income"

output=600+480=1080